OpenStudy (anonymous):

Identify the missing factors: X3 =x2 ( )

OpenStudy (anonymous):

x^3 = x^2 * y y = x^3 / x^2 y = x missing factor = x Please stop posting the same missing factor things. I already explained the "how to solve" process to you. If you don't understand please tell me

OpenStudy (anonymous):

what is it?